4. Find the zero of the polyrwmial in each of the following
cases:

U. p(x) = x + 5
(iv) p(x) = 3x - 2

(u) p(x) :::: x - 5 (iii) p(x) = 2x + 5
(u) p(x) = 3x (ui) p(x) = m:, a if 0

(uii) p(x) = e:t + d, c if 0, c, d are real numbers.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) For zero, p(x) = 0 => x + 6 = 0 => x = - 5 is a zero of the polynomial p(x). (il) For zero, p(x) = 0 => x - 5 = 0 x = 5 is a zero of the polynomial p(x ), (iii) For zero, p(x) • 0 => 2x + 5 = 0 => x = - £ is a zero of the polynomial p(x). (iv) For zero, p(x) = 0 => 3x - 2 = 0 => x = is a zero of the polynomial p(x). (v) For zero, p(x) = 0 => 3x = 0 => x = 0 is a zero of the polynomial p(x). (vi) For zero, p(x) = 0 => ax = 0 => x = 0, as a * 0 Therefore, x = 0 is a zero of the polynomial p(x). (uii) For zero, p(x) = 0 => ex + d = 0 d => x = - -, c Therefore, x = - d is a zero of the polynomial p). c